OpenStudy (anonymous):

Suppose x is a number such that 2x=x/2-6. Must it then be true that 3x=12?

OpenStudy (texaschic101):

2x = (x/2) - 6 --- multiply every term by 2 4x = x - 12 4x - x = 12 3x = 12 TRUE

OpenStudy (anonymous):

Thank you.

OpenStudy (texaschic101):

YW :)

OpenStudy (texaschic101):

wait.......stop.....mistake 2x = (x/2) - 6 -- multiply by 2 4x = x - 12 3x = -12 FALSE

OpenStudy (texaschic101):

sorry about that :(